OpenStudy (anonymous):

A principal of $450 is deposited in an account that pays 2.5% interest compounded yearly. Find the account balance after 3 years rounded to the nearest cent.

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) A = 450(1+0.025/1)^(1*3) A = ?? Tell me what you get

OpenStudy (anonymous):

450+11.25(3)

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do i need one

jimthompson5910 (jim_thompson5910):

yes you do, but you can use google as a calculator type "450(1+0.025/1)^(1*3)" without quotes into google and tell me what result you get

OpenStudy (anonymous):

i got a big number

jimthompson5910 (jim_thompson5910):

tell me what you got